OpenStudy (anonymous):

(3x+4 / 12) * (8 / 9x+12)

OpenStudy (anonymous):

(3x+4/3)*(2/9x+12) because 4*3=12 and 4*2=8 (3x+4/3)*{2/3(3x+4)} (1/3)*(2/3) (2/9)

OpenStudy (anonymous):

in 1st step 12 and 8 cancel out to 3 and 2

OpenStudy (anonymous):

in 3rd step 3x+4 cancels out with 3x+4